What is the present value of $7800.00 payable in six years if the current interest rate is 7.6% p.a., compounded quarterly? Round to nearest 100th.
Answer #1

Answer

  • Annual rate = 7.6%
  • No of quarter in a year = 4
  • Applicable rate = 7.6/4 = 1.90%
  • Period = 6 years
  • No. of quarters in 6 years = 6 x 4 = 24
  • Present value factor at 1.90% for 24th period = [ 1/(1.019)24] = 0.636531036
  • Present Value of $ 7800 = 7800 x 0.636531036 = $ 4964.94
  • ANSWER rounded to 100th = $ 5,000